Maria lights two candles of equal length at the same time. One candle takes six hours to burn out and the other takes nine. How much time will pass until the slower-burning candle is exactly twice as long as the faster-burning one?
Answer by scott8148(6628)

let x=time __ 2[(6-x)/6]=(9-x)/9 __ multiplying by 9 __ 3(6-x)=9-x __ 18-3x=9-x

adding 3x-9 __ 9=2x __ 4.5=x